OpenStudy (anonymous):

Find the volume of the solid obtained by rotating the region bounded by y = x and y = x^2 about the line x = -2.

OpenStudy (anonymous):

Those are the choices I was given as a result.. (5/6)pi (14/3)pi (Pi/6) (11/3)Pi

OpenStudy (zarkon):

\[\frac{5\pi}{6}\]

OpenStudy (zarkon):

\[2\pi\int\limits_{0}^{1}(x+2)(x-x^2)dx\]

OpenStudy (anonymous):

Ahhhhhhhhh...I knew I was forgetting something. calc 2 was so long ago :(

OpenStudy (zarkon):

you were close...just needed to make the radius bigger

OpenStudy (anonymous):

I knew I had to do something with that -2 >.> lol

OpenStudy (zarkon):

could have also done washers \[π\cdot\int\limits_{0}^{1}[(\sqrt{y}+2)^{2}-(y+2)^2]dy=\frac{5\pi}{6}\]